For a group of four men, the probability for the number, x, who live through the next year is given by

x

P(x)

0

0.045

1

0.015

2

0.1

3

0.55

4

0.29

(a) Find the mean and standard deviation of the number of men in a group of four who will live through the next year.
(b) Find the probability that at least 2 men in a group of four will survive the next year.

(c) Is 1 a significantly low number of men in a group of four to survive the next year? Why or why not?
Answer #1
x f(x) xP(x) x2P(x)
0 0.0450 0.00 0.0000
1 0.0150 0.02 0.0150
2 0.1000 0.20 0.4000
3 0.5500 1.65 4.9500
4 0.2900 1.16 4.6400
total 3.0250 10.0050
E(x) =μ= ΣxP(x) = 3.0250
E(x2) = Σx2P(x) = 10.0050
Var(x)=σ2 = E(x2)-(E(x))2= 0.8544
std deviation=         σ= √σ2 = 0.924

a) mean =3.025

standard deviation =0.924

b)

  probability that at least 2 men in a group of four will survive the next year=0.1+0.55+0.29 =0.94

c) No a sprobability of getting 1 or less is (0.045+0.015=0.06) is not less than 0.05 , therefore this is not an unusual event